OpenStudy (loser66):

Find the radius of \[\sum_{n=0}^\infty 3^n x^{n!}\] Please, help

geerky42 (geerky42):

Can you help? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

hmm interesting one here :)

OpenStudy (xapproachesinfinity):

hmm i'm thinking that logs would be useful here somehow lol don't know how yet

OpenStudy (anonymous):

Lets use the ratio test: \[\lim _{n \to \infty}\frac{3^{n+1}*|x|^{(n+1)!}}{3^n*|x|^{n!}}\] \[\frac{3^{n+1}*|x|^{(n+1)!}}{3^n*|x|^{n!}}=3*|x|^{(n+1)!-n!}=3*|x|^{n*n!}\]

OpenStudy (loser66):

the last one should be \(3x^{n+1}\) but then??

OpenStudy (xapproachesinfinity):

hmm there is an error in your simplification

OpenStudy (xapproachesinfinity):

good catch old man about to say that :)

OpenStudy (loser66):

@xapproachesinfinity old man has old eyes hahaha.. although it is slowly catching up but it works well, right? young boy??

OpenStudy (xapproachesinfinity):

hmm thinking how can that help

OpenStudy (xapproachesinfinity):

we take the limit and keep track of possible conditions of x lol

OpenStudy (xapproachesinfinity):

3|x|.|x|^n limit of x^n?

OpenStudy (xapproachesinfinity):

since we have the || that gotta either be 0 or infinity

OpenStudy (xapproachesinfinity):

can we have cases lol

OpenStudy (loser66):

Forget it. I just wonder how to work on it. It is not my real problem

OpenStudy (anonymous):

\[\frac{3^{n+1}|x|^{(n+1)!}}{3^n|x|^{n!}}=3|x|^{(n+1)!-n!}=3|x|^{(n+1)*n!-n!}=3|x|^{n!((n+1)-1)}=3|x|^{n!*n}\]

OpenStudy (xapproachesinfinity):

wolfram said it converges where did you find the problem

OpenStudy (xapproachesinfinity):

eh that is actually right :)

OpenStudy (xapproachesinfinity):

but it doesn't help to get anywhere

OpenStudy (xapproachesinfinity):

let c be the finite number that is converges to \[c=\sum 3^nx^{n!} \Longrightarrow \ln c=\sum \ln (3^nx^{n!}) \] \[=\sum (n\ln 3+x\ln(n!))\]

OpenStudy (xapproachesinfinity):

some crazy nonsense here hehe

OpenStudy (xapproachesinfinity):

\[=\ln3\sum n+x\sum\ln n!\] sum n diverges lol

OpenStudy (xapproachesinfinity):

@Loser66

OpenStudy (xapproachesinfinity):

you give up haha

geerky42 (geerky42):

So what you are saying is that this sum never converges regardless the value of x, right?

OpenStudy (xapproachesinfinity):

oh no! something is wrong there! i don't even know if that's allowed to do. the series converge when i checked it in wolfram

OpenStudy (xapproachesinfinity):

@Zarkon

OpenStudy (xapproachesinfinity):

we need some insight :)

OpenStudy (xapproachesinfinity):

eh made a mistake \[=\ln3\sum n+\ln x\sum n!\]

OpenStudy (xapproachesinfinity):

now both n! and n are divergent series! lol that means this method won't give anything useful

OpenStudy (anonymous):

Why cant we use the ratio test and then use diff cases for x?

OpenStudy (anonymous):

So when |x| is less than 1 then the series converges to 0 If |x| equals 1 then the series converges to 3*1=3 And if |x| is greater than 1 then it diverges and goes to \(\infty\)

OpenStudy (anonymous):

whoops n*n! goes to infinity as n goes to infinty

OpenStudy (xapproachesinfinity):

it does not converges o the limit is zero but the series not necessary to 0 for the first case you mentioned

OpenStudy (xapproachesinfinity):

but i think this series is beyond just a ratio test there is a big subtlety with that limit

OpenStudy (xapproachesinfinity):

actually when the limit is zero the radius is infinite

OpenStudy (xapproachesinfinity):

in other words converges for any x

OpenStudy (xapproachesinfinity):

@SithsAndGiggles

OpenStudy (anonymous):

The root test only works for series of the form \(\sum a_n x^n\), not \(\sum a_n x^{n!}\).

OpenStudy (xapproachesinfinity):

what about ratio that was ratio test not root

OpenStudy (anonymous):

I actually meant "ratio" :P but yes the same goes for ratio test.

OpenStudy (xapproachesinfinity):

i guess ratio won't work as well

OpenStudy (xapproachesinfinity):

how can we do this one

OpenStudy (anonymous):

@Loser66 had a very similar problem before, but then we had \(x^{n^2}\). Setting \(k=n^2\) worked, but I don't think a similar substitution will work this time around.

OpenStudy (xapproachesinfinity):

what if we you some N! apprax'ts

OpenStudy (xapproachesinfinity):

use*

OpenStudy (anonymous):

We might be able to make a fair guess about the interval of convergence, though. Clearly the series diverges if \(x=\pm1\), since that gives us a geometric series \(\sum 3^n\). I think it might be \(|x|<1\), so the radius would be \(1\).

OpenStudy (anonymous):

I like the approximation suggestion. Stirling's approximation says \[n!\approx \sqrt{2\pi n}\left(\frac{n}{e}\right)^n\] I'm not sure if this will get us anywhere though. There might be some trick we can use involving logarithms.

OpenStudy (loser66):

That is my problem :) I don't know how to turn it to the form \(a_kx^k\)

OpenStudy (xapproachesinfinity):

that needs some real work

OpenStudy (xapproachesinfinity):

gotta leave :)

OpenStudy (loser66):

@SithsAndGiggles How about this problem? Find the radius of convergent series \[\sum_{n=0}^\infty \dfrac{5^n}{n!}x^n\]

OpenStudy (loser66):

My other question on it is : Is there any difference if n =1 to infinitive instead of n =0?

OpenStudy (anonymous):

The ratio test works quite nicely. \[\lim_{n\to\infty}\left|\frac{\dfrac{5^{n+1}}{(n+1)!}x^{n+1}}{\dfrac{5^n}{n!}x^n}\right|=5|x|\lim_{n\to\infty}\frac{n!}{(n+1)!}\] The starting index of the series does not matter.

OpenStudy (loser66):

So, the radius =\(\infty\) , right?

OpenStudy (anonymous):

Right, since \(0<1\), the second series converges everywhere.

OpenStudy (loser66):

How about this: I would like to apply Hadamard test this time, please \[\sum_{n=0}^\infty \dfrac{n^n}{n!}x^n\]

OpenStudy (loser66):

The net is so bad. :(

OpenStudy (anonymous):

Let's use a strategy similar to the one you had before. \[a_k=\begin{cases}3^n&\text{for }k=n!\\ 0&\text{otherwise}\end{cases}\] Then \[\sqrt[k]{a_k}=\begin{cases}\large \sqrt[n!]{3^n}&\text{for }k=n!\\ 0&\text{otherwise}\end{cases}=\begin{cases}3^{1/(n-1)!}&\text{for }k=n!\\ 0&\text{otherwise}\end{cases}\] Clearly, \(\limsup\limits_{k\to\infty}\sqrt[k]{a_k}=1\).